OpenStudy (anonymous):

find the point of intersection: x = 17+7t y = 10+5t z = 13+3t x = -20+8s y = -19+7s z = -8+6s

OpenStudy (anonymous):

i solved the first equation for s...solved the 2nd equation for t...then sub'd both into the 3rd equation

OpenStudy (anonymous):

and now i'm stuck

OpenStudy (anonymous):

i got t = 3 and s = 5

jimthompson5910 (jim_thompson5910):

if t = 3, then x = 17+7t = 17+7(3) = 17+21 = 38 but if s = 5, then x = -20+8s = -20+8(5) = -20+40 = 20 So the two x values aren't equal, which means that the solution for (s,t) isn't (5,3)

OpenStudy (anonymous):

yea...i've gathered

jimthompson5910 (jim_thompson5910):

the system x = 17+7t y = 10+5t z = 13+3t x = -20+8s y = -19+7s z = -8+6s can be stated as -20+8s = 17+7t -19+7s = 10+5t -8+6s= 13+3t which can then be rewritten as 8s-7t = 37 7s-5t = 29 6s-3t = 21 Did you get this for your work?

OpenStudy (anonymous):

no..where did u get the 37,29 and 21?

OpenStudy (anonymous):

oh okay i see

jimthompson5910 (jim_thompson5910):

I moved all the constant terms to the right side

jimthompson5910 (jim_thompson5910):

solve that system to find your solution (s,t)